OpenStudy (anonymous):

Sqrt 4xy^3 * sqrt 6y

OpenStudy (xapproachesinfinity):

\[\text{i suppose the question is to simplify?!}\]

OpenStudy (anonymous):

Yes it's simplify sorry

OpenStudy (xapproachesinfinity):

\[\sqrt{4xy^3}\times \sqrt{6y}\] first use one property of radicals \[\sqrt{4xy^3}\times \sqrt{6y}=\sqrt{4xy^36y}\]

OpenStudy (xapproachesinfinity):

this the property \[\sqrt{a}\sqrt{b}=\sqrt{ab}\]

OpenStudy (xapproachesinfinity):

now we have \[=\sqrt{4xy^46}=2y\sqrt{6x}\]

OpenStudy (xapproachesinfinity):

can you tell me where 2 came from and also y?

OpenStudy (anonymous):

Square root of 4=2

OpenStudy (xapproachesinfinity):

yes that's correct

OpenStudy (anonymous):

4 is a perf square so it can be simplified

OpenStudy (xapproachesinfinity):

and also \[\sqrt{y^4} =y^2\] y^4 is perfect square too

OpenStudy (xapproachesinfinity):

i realized i made a mistake above should have gotten \[2y^2\sqrt{6x}\]

OpenStudy (anonymous):

Thank you can you stay and help? I'm taking online classes and don't have much help